# Strong semistability of Higgs bundles over curves

### v4

@@ Abstract.
In this paper we complete the study of the Lan-Sheng-Zuo conjecture proposed in [LSZ13] for the curve case. Precisely, we prove that every semistable Higgs bundle is strongly semistable for curves of genus $g\leq 1$, and over any curves of genus $g\geq 2$ construct explicit examples of semistable Higgs bundles of arbitrary big rank (the first example is $p=2,r=3$) which are not strongly semistable. These results are complementary to the strongly semistability theorem of Lan-Sheng-Yang-Zuo [LSZ19] and Langer [Lan14] for semistable Higgs bundles of small rank.
@@ 1. Introduction
In this paper we work over an algebraically closed field $k$ with $\text{char}(k)=p>0$, and curves mean nonsingular projective curves over $k$.
Let $C$ be a curve and $\mathcal{V}$ be a vector bundle over $C$. Classically, $\mathcal{V}$ is called strongly semistable if for all $n\geq 0$, $(F^{n})^{*}\mathcal{V}$ is semistable, where $F$ is the absolute Frobenius of $C$. It is natural to ask if every semistable vector bundle is strongly semistable. An answer in either direction is interesting to know, since $F_{C}$ is inseparable (pullback along a seperable morphism preserves semistability). D. Gieseker [Gie73] constructed a rank two semistable vector bundle over curve which is not strongly semistable.
Based on the advance [OV07] in non-abelian Hodge theory in positive characteristic, Lan-Sheng-Zuo [LSZ13] introduced the notion of strongly semistability of Higgs bundles, and proved that every rank two nilpotent semistable Higgs bundle is strongly semistable ([LSZ13, Theorem 2.6]). In particular, the example of Gieseker is actually strongly semistable in this new sense, regarded as a Higgs bundle with zero Higgs field. Then it becomes another natural question that whether or not every nilpotent  [^1] The nilpotency condition is a necessary condition for a Higgs bundle to initialize a Higgs-de Rham flow. semistable Higgs bundle is strongly semistable. With an optimistic attitude, Lan-Sheng-Zuo proposed the following conjecture:
@@
Every semistable nilpotent Higgs bundle of exponent $\leq p-1$ is strongly semistable.
In ([Li14, Theorem 1.1]), Li proved the conjecture for the rank three case over curves under the assumption $\text{char}(k)\geq 3$. Accepted as a common terminology nowadays, the rank of a vector bundle is said to be small, when it is $\leq p$, and hence will be called big in this paper, if the otherwise holds. Hence it remains undetermined whether Li’s result can be extended to the first big rank case, namely $p=2,r=3$. On the other hand, the next result extends Li’s result to an arbitrary small rank, which shall be called the strongly semistability theorem.
Theorem 1.1 ([LSZ19, Theorem A.1][Lan14, Theorem 5.5]).
@@
Let $p$ be any prime number and $C$ be a smooth projective curve over $k$ with $\text{char}(k)=p$. The following two statements hold:
  1. (i)
@@
  2. (ii)
    If $g(C)\geq 2$, then for any big rank $r$, there exists a rank $r$ semistable Higgs bundle over $C$ which is not strongly semistable;
  3. (iii)
    If $C$ is generic of $g(C)\geq 2$, then the category consisting of strongly semistable Higgs bundles over $C$ is not closed under tensor product.
Note that for $C=\mathbb{P}^{1}$, Conjecture 1.1 holds trivially, since every semistable Higgs bundle on $\mathbb{P}^{1}$ is a direct sum of line bundles of the same degree, equipped with the zero Higgs field. To show $(iii)$, we construct two strongly semistable Higgs bundles whose tensor product is semistable but not strongly semistable. For $p=2$, the construction actually extends to every curve. We believe it is also the case when $p$ is odd. This result negates our earlier hope, although the category of semistable Higgs bundles over $C$ does not form a tensor category, the subcategory consisting of strongly semistable Higgs bundles might enjoy this property, making a stronger analogue to the category of semistable Higgs bundles in char zero.
The fundamental theorem of Deligne-Illusie in Hodge theory in positive characteristic asserts that, for a $W_{2}(k)$-liftable smooth proper $k$-variety of dimension $\leq p$, the Hodge to de Rham spectral sequence degenerates at $E_{1}$-page. In a very recent work by A. Petrov [Pet23, Theorem 1.1], he constructed a $W_{2}(k)$-liftable smooth proper $k$-variety of dimension $p+1$ such that the $E_{1}$-degeneration property fails. Our result may be viewed as an analogue in non-abelian Hodge theory in positive characteristic. Let us notice that once a counterexample in the case $r=p+1$ is constructed, a counterexample of every big rank follows immediately (see Proposition 4.1, Corollary 4.1).
@@ Acknowledgment
The authors would like to thank Jianping Wang for his comments on Remark 3.1, which improves previous result, and to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments which improved the paper. The second named author is supported by the Chinese Academy of Sciences Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@ 1. $(1)$
    We say that $(\mathcal{E},\theta)$ is nilpotent of exponent $\ell$ if for any open subset $U\subset X$ and any collection $\left\{\partial_{x_{1}},\ldots,\partial_{x_{\ell}}\right\}$ of sections of $\mathcal{T}_{X}(U)$, we have
    $\theta\left(\partial_{x_{1}}\right)\ldots\theta\left(\partial_{x_{\ell}}\right)=0.$
@@ 2. $(2)$
    There exists a filtration of saturated Higgs subsheaves
    $0=\mathcal{F}^{m}\subset\mathcal{F}^{m-1}\subset\ldots\subset\mathcal{F}^{0}=(\mathcal{E},\theta)$
    of length $m\leq\ell$ such that the associated graded Higgs bundle has zero Higgs field.
@@
Let $(\mathcal{E},\theta)$ be a Higgs bundle on $X$. A Higgs-de Rham flow with initial term $(\mathcal{E},\theta)$ is an infinite sequence
@@
in which $C^{-1}$ is the inverse Cartier transform with respect to a given $W_{2}(k)$-lifting $X\hookrightarrow\widetilde{X}$ and $\mathcal{N}_{i}^{\bullet}$ is a Griffiths transverse filtration of $(\mathcal{V}_{i},\nabla_{i})$ such that $(\mathcal{E}_{i+1},\theta_{i+1})$ is the associated graded Higgs bundle for each $i$.
@@
where $\mathcal{N}_{i}^{\bullet}$ is the trivial filtration for each $i$. This shows the notion of strongly semistability for a Higgs bundle generalizes naturally the notion of strongly semistability for a vector bundle.
@@
The following existence theorem of gr-semistable filtration on semistable flat bundles is the key to the strongly semistability theorem (Theorem 1.1).
Theorem 2.2 ([LSZ19, Theorem A.4][Lan14, Theorem 5.5]).
@@
In general, gr-semistable filtration on a semistable flat bundle is not unique, but the associated graded Higgs bundles with respect to different gr-semistable filtrations are $S$-equivalent.
Lemma 2.2 ([Lan14, Corollary 5.6]).
If $\mathcal{N}_{1}^{\bullet}$ and $\mathcal{N}_{2}^{\bullet}$ are gr-semistable filtrations of a $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$, then the associated graded Higgs bundles are $S$-equivalent. As a consequence, if the associated graded Higgs bundle is stable, then the gr-semistable filtration is unique.
@@
Let $(\mathcal{V},\nabla)$ be a semistable flat bundle. If there exists a gr-polystable filtration, then the exponent of nilpotency of the associated graded Higgs bundle with respect to gr-polystable filtration is less or equal to the exponent of nilpotency of the associated graded Higgs bundle with respect to any gr-semistable filtration.
@@
Since $(\mathcal{E},\theta)$ is not polystable, there exists a non-split exact sequence of Higgs bundles
$0\to(\mathcal{E}^{\prime},\theta^{\prime})\to(\mathcal{E},\theta)\to(\mathcal{E}^{\prime\prime},\theta^{\prime\prime})\to 0,$
and the maximum between the exponents of nilpotency of $(\mathcal{E}^{\prime},\theta^{\prime})$ and $(\mathcal{E}^{\prime\prime},\theta^{\prime\prime})$ is less or equal to the exponent of nilpotency of $(\mathcal{E},\theta)$. If both $(\mathcal{E}^{\prime},\theta^{\prime})$ and $(\mathcal{E}^{\prime\prime},\theta^{\prime\prime})$ are polystable, then $(\widetilde{\mathcal{E}},\widetilde{\theta})\cong(\mathcal{E}^{\prime},\theta^{\prime})\oplus(\mathcal{E}^{\prime\prime},\theta^{\prime\prime})$ and this completes the proof. Otherwise we repeat the same argument to $(\mathcal{E}^{\prime},\theta^{\prime})$ and $(\mathcal{E}^{\prime\prime},\theta^{\prime\prime})$, after finite steps this also completes the proof. ∎
@@
Let $E$ be an elliptic curve over $k$ with $\text{char}(k)=2$. Let $\mathcal{E}$ be the non-trivial extension of structure sheaf $\mathcal{O}_{E}$ by a degree one line bundle. Then $\mathcal{E}$ is stable while $F^{*}\mathcal{E}$ is semistable but not stable, since $F^{*}\mathcal{E}$ is of degree two and rank two. Suppose a Jordan-Hölder filtration of $F^{*}\mathcal{E}$ is given by
@@
The only possible gr-semistable filtrations of $(F^{*}\mathcal{E},\nabla_{\operatorname{can}})$ are given by its Jordan-Hölder filtration or the trivial filtration. By [Oda71, Theorem 2.16] we know that $F^{*}\mathcal{E}$ is indecomposable. On the other hand, $\mathcal{E}_{1}$ is not $\nabla_{\operatorname{can}}$-invariant. This shows that in any case the associated graded Higgs bundle is not polystable.
@@
[…]
@@ 3. Lan-Sheng-Zuo conjecture on elliptic curves
be the Harder-Narasimhan filtration of $\mathcal{V}$. Then $\mathcal{V}^{s}$ is not $\nabla$-invariant since $(\mathcal{V},\nabla)$ is $\nabla$-semistable. Then there is a unique $\ell<s$ such that $\nabla(\mathcal{V}^{s})\subsetneq\mathcal{V}^{\ell+1}$ but $\nabla(\mathcal{V}^{s})\subseteq\mathcal{V}^{\ell}$. Then $\nabla$ induces a non-zero morphism $\varphi\colon\mathcal{V}^{s}\to\mathcal{G}^{\ell}:=\mathcal{V}^{\ell}/\mathcal{V}^{\ell+1}$, which is a contradiction, as $\mu(\mathcal{V}^{s})>\mu(\mathcal{G}^{\ell})$.
Since $\mathcal{V}$ is a semistable bundle, we may choose the trivial filtration on $(\mathcal{V},\nabla)$ as its gr-semistable filtration, and the associated graded Higgs bundle is $(\mathcal{V},0)$. If $\mathcal{V}$ is a semistable bundle, then $(F^{*}\mathcal{V},\nabla_{\operatorname{can}})$ is $\nabla_{\operatorname{can}}$-semistable, and by above claim we have $F^{*}\mathcal{V}$ is a semistable bundle. This shows every semistable bundle on elliptic curve is strongly semistable.

### v5

@@ Abstract.
In this paper we complete the study of the Lan-Sheng-Zuo conjecture proposed in [LSZ13] for the curve case. Precisely, we prove that every nilpotent semistable Higgs bundle of exponent $\leq p-1$ is strongly semistable for curves of genus $g\leq 1$, and over any curves of genus $g\geq 2$ construct explicit examples of nilpotent semistable Higgs bundles of exponent $\leq p-1$ and arbitrary big rank (the first example is $p=2$ and rank $3$) which are not strongly semistable. These results are complementary to the strong semistability theorem of Lan-Sheng-Yang-Zuo [LSZ19] and Langer [Lan14] for semistable Higgs bundles of small rank.
@@ 1. Introduction
In this paper we work over an algebraically closed field $k$ with $\mathrm{char}(k)=p>0$, and curves mean nonsingular projective curves over $k$.
Let $C$ be a curve and $\mathcal{V}$ be a vector bundle over $C$. Classically, $\mathcal{V}$ is called strongly semistable if for all $n\geq 0$, $(F^{n})^{*}\mathcal{V}$ is semistable, where $F$ is the absolute Frobenius of $C$. It is natural to ask if every semistable vector bundle is strongly semistable. An answer in either direction is interesting to know, since $F$ is inseparable (pullback along a finite seperable morphism preserves semistability). D. Gieseker [Gie73] constructed a rank two semistable vector bundle over curve which is not strongly semistable.
Let $X$ be a smooth projective variety defined over $k$ which is $W_{2}(k)$-liftable. Based on the advance [OV07] in non-abelian Hodge theory in positive characteristic, Lan-Sheng-Zuo [LSZ13] introduced the notion of strong semistability of Higgs bundles on $X$, and proved that every rank two nilpotent semistable Higgs bundle is strongly semistable ([LSZ13, Theorem 2.6]). In particular, the example of Gieseker is actually strongly semistable in this new sense, regarded as a Higgs bundle with zero Higgs field. Then it becomes another natural question that whether or not every nilpotent  [^1] The nilpotency condition is a necessary condition for a Higgs bundle to initialize a Higgs-de Rham flow. semistable Higgs bundle of exponent $\leq p-1$ is strongly semistable. With an optimistic attitude, Lan-Sheng-Zuo proposed the following conjecture:
@@
Every nilpotent semistable Higgs bundle of exponent $\leq p-1$ is strongly semistable.
In ([Li14, Theorem 1.1]), Li proved the conjecture for the rank three case over curves under the assumption $\mathrm{char}(k)\geq 3$. Accepted as a common terminology nowadays, the rank of a vector bundle is said to be small, when it is $\leq p$, and hence will be called big in this paper, if the otherwise holds. Hence it remains undetermined whether Li’s result can be extended to the first big rank case, namely $p=2$ and rank $3$. On the other hand, the next result extends Li’s result to an arbitrary small rank, which shall be called the strong semistability theorem.
Theorem 1.1 ([LSZ19, Theorem A.1], [Lan14, Theorem 5.5]).
@@
Let $p$ be any prime number and $C$ be a smooth projective curve over $k$ with $\mathrm{char}(k)=p$. The following three statements hold:
  1. (1)
@@
  2. (2)
    If $g(C)\geq 2$, then for any big rank $r$, there exists a rank $r$ nilpotent semistable Higgs bundle of exponent $\leq 1$ over $C$ which is not strongly semistable;
  3. (3)
    Suppose $p$ is odd and $C$ is generic of $g(C)\geq 2$, or $p=2$ and $g(C)\geq 2$. Then there exist a nilpotent strongly semistable Higgs bundles of exponent $\leq 1$ and that of exponent =0=0 such that their tensor product is not strongly semistable.
Note that for $C=\mathbb{P}^{1}$, Conjecture 1.1 holds trivially, since every semistable Higgs bundle on $\mathbb{P}^{1}$ is a direct sum of line bundles of the same degree, equipped with the zero Higgs field. We believe $(3)$ remains true when $p$ is odd without assuming that $C$ is generic. This result negates our earlier hope, although the category of semistable Higgs bundles over $C$ does not form a tensor category, the subcategory consisting of strongly semistable Higgs bundles might enjoy this property, making a stronger analogue to the category of semistable Higgs bundles in characteristic zero.
The fundamental theorem of Deligne-Illusie in Hodge theory in positive characteristic asserts that, for a $W_{2}(k)$-liftable smooth proper $k$-variety of dimension $\leq p$, the Hodge to de Rham spectral sequence degenerates at $E_{1}$-page. In a very recent work by A. Petrov [Pet23, Theorem 1.1], he constructed a $W_{2}(k)$-liftable smooth proper $k$-variety of dimension $p+1$ such that the $E_{1}$-degeneration property fails. Our result may be viewed as an analogue in non-abelian Hodge theory in positive characteristic.
@@ Acknowledgment
The authors would like to thank Jianping Wang for his comments on Remark 3.1, which improves previous result, and to thank Lingguang Li, for his careful reading of the manuscript and for providing valuable comments which improved the paper. They would like to thank heartily an anonymous referee for scrutinizing the work and providing valuable comments and suggestions, which have greatly improved the quality of the paper. The second named author is supported by the Chinese Academy of Sciences Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@ 2. Preliminaries on Higgs-de Rham flows
    We say that $(\mathcal{E},\theta)$ is nilpotent of exponent $\ell$ if for any open subset $U\subset X$ and any collection $\left\{\partial_{x_{1}},\ldots,\partial_{x_{\ell}}\right\}$ of sections of $\mathcal{T}_{X}(U)$, we have θ(∂_x_1) …θ(∂_x_ℓ)=0.
@@
  3. (3)
    We say that $(\mathcal{E},\theta)$ is nilpotent of exponent $\leq\ell$ if it is either nilpotent of exponent $\ell$ or $\ell$.
@@
    There exists a filtration of saturated Higgs subsheaves 0=F^m F^m-1 F^0=(E, θ) of length $m\leq\ell$ such that the associated graded Higgs bundle has zero Higgs field.
@@
Let $(\mathcal{E},\theta)$ be a nilpotent Higgs bundle on $X$ of exponent $\leq p-1$. A Higgs-de Rham flow with initial term $(\mathcal{E},\theta)$ is an infinite sequence
@@
in which $C^{-1}$ is the inverse Cartier transform with respect to a given $W_{2}(k)$-lifting $X\hookrightarrow\widetilde{X}$ and $\mathcal{N}_{i}^{\bullet}$ is a Griffiths transverse filtration of $(\mathcal{V}_{i},\nabla_{i})$ such that $(\mathcal{E}_{i+1},\theta_{i+1})$ is the associated graded Higgs bundle for each $i$ which is nilpotent of exponent $\leq p-1$.
@@
where $\mathcal{N}_{i}^{\bullet}$ is the trivial filtration for each $i$. This shows the notion of strong semistability for a Higgs bundle generalizes naturally the notion of strong semistability for a vector bundle.
From now on, we will omit to mention the fixed ample line bindle $\omega$ when we consider (semi)stability.
@@
The following existence theorem of gr-semistable filtration on semistable flat bundles is the key to the strong semistability theorem (Theorem 1.1).
Theorem 2.2 ([LSZ19, Theorem A.4], [Lan14, Theorem 5.5]).
@@
In general, gr-semistable filtration on a semistable flat bundle is not unique, but associated graded Higgs bundles with respect to different gr-semistable filtrations are $S$-equivalent.
Lemma 2.2 ([LSZ19, Proposition 6.9], [Lan14, Corollary 5.6]).
If $\mathcal{N}_{1}^{\bullet}$ and $\mathcal{N}_{2}^{\bullet}$ are gr-semistable filtrations of a $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$, then the reflexivizations of the associated graded slope polystable Higgs bundles obtained from the Jordan-Hölder filtrations of the associated graded semistable Higgs bundles are isomorphic. Moreover, if the associated graded Higgs bundle is stable, then the gr-semistable filtration is unique.
@@
Let $(\mathcal{V},\nabla)$ be a semistable flat bundle. If there exists a gr-polystable filtration, then the exponent of nilpotency of the associated graded Higgs bundle with respect to gr-polystable filtration is less than or equal to the exponent of nilpotency of the associated graded Higgs bundle with respect to any gr-semistable filtration.
@@
$0=(\mathcal{E}_{0},\theta_{0})\subset(\mathcal{E}_{1},\theta_{1})\subset\dots\subset(\mathcal{E}_{m},\theta_{m})=(\mathcal{E},\theta)$
be a Jordan-Hölder filtration of $(\mathcal{E},\theta)$ and let $(\overline{\mathcal{E}}_{i},\overline{\theta}_{i})=(\mathcal{E}_{i},\theta_{i})/(\mathcal{E}_{i-1},\theta_{i-1})$ for $1\leq i\leq m$. Then the maximum among the exponents of nilpotency $(\overline{\mathcal{E}}_{i},\overline{\theta}_{i})$ is less than for equal to the exponent of nilpotency of $(\mathcal{E},\theta)$.
Moreover, $(\widetilde{\mathcal{E}},\widetilde{\theta})^{**}\cong\bigoplus_{i=1}^{m}(\overline{\mathcal{E}}_{i},\overline{\theta}_{i})^{**}$ (where ** denotes the reflexivization) by Lemma 2.2, and this completes the proof because the reflexivization does not change the exponent of nilpotency. ∎
@@
Let $E$ be an elliptic curve over $k$ with $\mathrm{char}(k)=2$. Let $\mathcal{E}$ be the non-trivial extension of structure sheaf $\mathcal{O}_{E}$ by a degree one line bundle. Then $\mathcal{E}$ is stable while $F^{*}\mathcal{E}$ is semistable, as every semistable bundle on elliptic curve is strongly semistable. However, $F^{*}\mathcal{E}$ is not stable, since there is no stable bundle of rank $2$ and degree $2$ on elliptic curve.
Suppose a Jordan-Hölder filtration of $F^{*}\mathcal{E}$ is given by
@@
The only possible gr-semistable filtrations of $(F^{*}\mathcal{E},\nabla_{\operatorname{can}})$ are given by its Jordan-Hölder filtration or the trivial filtration. We know that $F^{*}\mathcal{E}$ is indecomposable by [Oda71, Theorem 2.16] and that $\mathcal{E}_{1}$ is not $\nabla_{\operatorname{can}}$-invariant. This shows that in any case the associated graded Higgs bundle is not polystable.
@@
Let $C$ be a genus $g\geq 2$ curve and $\mathcal{V}$ be a non-trivial extension of $\mathcal{O}_{C}$ by $\mathcal{O}_{C}$. Since $\mathcal{V}$ is an indecomposable bundle of degree zero, there exists a connection $\nabla$ on it by [BS06, Proposition 3.1]. Without loss of generality, we may assume that $\nabla$ does not preserve the subbundle $\mathcal{O}_{C}$.
Indeed, consider the sheaf $\mathcal{A}$ of endomorphisms $\varphi\colon\mathcal{V}\to\mathcal{V}\otimes K_{C}$ which maps the subbundle $\mathcal{O}_{C}$ into $\mathcal{O}_{C}\otimes K_{C}$. Then $\varphi\mapsto\varphi|_{\mathcal{O}_{C}}$ gives a map $\mathcal{A}\to K_{C}$ and its kernel is $\mathcal{V}\otimes K_{C}$. Hence
$\dim H^{0}(C,\mathcal{A})\leq\dim H^{0}(C,K_{C})+\dim H^{0}(C,\mathcal{V}\otimes K_{C}).$
Since $\mathcal{V}$ is a non-trivial extension of $\mathcal{O}_{C}$ by $\mathcal{O}_{C}$, $\dim H^{0}(C,\mathcal{V}\otimes K_{C})=2g-1$ and so $\dim H^{0}(C,\mathcal{A})\leq 3g-1$.
On the other hand, since $\mathcal{V}$ is a non-trivial extension of $\mathcal{O}_{C}$ by $\mathcal{O}_{C}$, we see that $\dim H^{0}(C,\operatorname{\mathscr{E}\text{\kern-3.0pt{\large nd}}\,\,\,}(\mathcal{V}))=2$, and this fact, Serre duality and Riemann-Roch imply that $\dim H^{0}(\operatorname{\mathscr{E}\text{\kern-3.0pt{\large nd}}\,\,\,}(\mathcal{V})\otimes K_{C})=4g-2>3g-1$. So we may take an element $\xi\in H^{0}(C,\operatorname{\mathscr{E}\text{\kern-3.0pt{\large nd}}\,\,\,}\,(\mathcal{V})\otimes K_{C})$ which does not preserve the subbundle $\mathcal{O}_{C}$. Then $\nabla$ or $\nabla+\xi$ is such a connection as requested.
Claim that there is no gr-polystable filtration on $\mathcal{V}$. Indeed, note that
@@ 3. Lan-Sheng-Zuo conjecture on elliptic curves
be the Harder-Narasimhan filtration of $\mathcal{V}$. Then $\mathcal{V}^{s}$ is not $\nabla$-invariant since $(\mathcal{V},\nabla)$ is $\nabla$-semistable. Then there is a unique $\ell<s$ such that $\nabla(\mathcal{V}^{s})\not\subseteq\mathcal{V}^{\ell+1}$ but $\nabla(\mathcal{V}^{s})\subseteq\mathcal{V}^{\ell}$. Then $\nabla$ induces a non-zero morphism $\varphi\colon\mathcal{V}^{s}\to\mathcal{G}^{\ell}:=\mathcal{V}^{\ell}/\mathcal{V}^{\ell+1}$, which is a contradiction, as $\mu(\mathcal{V}^{s})>\mu(\mathcal{G}^{\ell})$.
Since $\mathcal{V}$ is a semistable bundle, we may choose the trivial filtration on $(\mathcal{V},\nabla)$ as its gr-semistable filtration, and the associated graded Higgs bundle is $(\mathcal{V},0)$. If $\mathcal{V}$ is a semistable bundle, then $(F^{*}\mathcal{V},\nabla_{\operatorname{can}})$ is $\nabla$-semistable, and by above claim we have $F^{*}\mathcal{V}$ is a semistable bundle. This shows every semistable bundle on elliptic curve is strongly semistable.